OpenStudy (anonymous):

Find the probability: A bag contains 4 red marbles, 16 yellow marbles, 5 purple marbles, 16 blue marbles, and 10 green marbles. What is the probability of pulling out a red or a green marble?

OpenStudy (kropot72):

P(red) = 4/51 P(green) = 10/51 The event red or green can happen in 2 different and mutually exclusive ways. So the probability of it happening is the sum of the separate probabilities. Can you now solve it?

OpenStudy (anonymous):

Would that be 14/51?

OpenStudy (kropot72):

Good work! You've got it :)

OpenStudy (anonymous):

That's the answer?

OpenStudy (kropot72):

Yes